OpenStudy (anonymous):

Simplify 13x - 4[x + (3 - x)]. 13x - 12 9x - 1 8x - 12

OpenStudy (anonymous):

what u got till now ?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

13x - 4[x + (3 - x)]. 13x-4(x+3-x)=13x-4(3)=13x-12

OpenStudy (anonymous):

i got a 150 question test an a hangover i got nothing

OpenStudy (anonymous):

i c good luck with the rest :P

OpenStudy (anonymous):

ouch

OpenStudy (anonymous):

150 questions will give anyone a hangover

OpenStudy (anonymous):

true

OpenStudy (anonymous):

you can probably do one a minute with this http://www.wolframalpha.com/input/?i=13x+-+4 [x+%2B+%283+-+x%29]

OpenStudy (anonymous):

let me try that again http://www.wolframalpha.com/input/?i=13x+-+4 [x+%2B+%283+-+x%29]

OpenStudy (anonymous):

ok copy and paste it gives \(13x-12\)

OpenStudy (anonymous):

wanna try another

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

7 - (2x + 11) + 3(3 - x) = 20 whats x epual